If the supply function in a market of face mask is Q = -10 + 10P and the demand function is 18 - 4P, and the government puts in a price floor at Php 3, what will the result be?


1
Expert's answer
2021-11-19T11:12:20-0500

In equilibrium Qd = Qs, so:

"18 - 4P = -10 + 10P,"

14P = 28,

P = 2,

Q = 18 - 4×2 = 10 units.

If the government puts in a price floor at Php 3, then there will be a surplus of:

"-10 + 10\u00d73 - (18 - 4\u00d73) = 14" units.